# Score-based Generative Priors Guided Model-driven Network for MRI Reconstruction

### v1

@@ VII Conclusions
In this study, we proposed a novel workflow to eliminate the need for optimizing the forward and reverse process in the SMLD-based method for MRI reconstruction. Instead, an unrolled model-driven network is proposed to extract valuable information from naive sampling results. Subsequently, the SMLD results can be periodically updated and utilized to guide the network training process. The experimental results demonstrated that the proposed hybrid method outperforms the workflows solely based on diffusion models or unrolled networks. Additionally, our method alleviates the substantial requirements for sampling steps and training images. When compared with cutting-edge methods, our network consistently produces reconstructions of higher quality, showcasing its superiority in MRI reconstruction tasks.

### v2

@@ 7 Conclusions
In this study, we proposed a novel MRI reconstruction workflow consisting of sampling, denoising and guided reconstruction steps. Naive SMLD samples are denoised in DMs and the artifacts are effectively mitigated. Then, the denoised samples are periodically updated and utilized to guide the network training process to recover fine details. Experimental results demonstrated that the proposed method outperforms the workflows solely based on diffusion models or unrolled networks. Additionally, our method alleviates the substantial requirements for sampling steps and training images. When compared with cutting-edge methods, our network consistently produces reconstructions of higher quality, showcasing its superiority in MRI reconstruction tasks.
8 Acknowledgements
This work was supported by the National Natural Science Foundation of China [Nos. 62331008, 62027827, 62221005 and 62276040], National Science Foundation of Chongqing [Nos. 2023NSCQ-LZX0045 and CSTB2022NSCQ-MSX0436], Chongqing University of Posts and Telecommunications ph.D. Innovative Talents Project [BYJS202211].